OpenStudy (mackenzie2013):

Solve for x. x + 47 = -23 x =

OpenStudy (anonymous):

subtract 47 from both sides

OpenStudy (anonymous):

x=-23-47 x=-70

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

hey can someone help me with a question i made please

OpenStudy (the_fizicx99):

X + 47 = -23 -47 -47 X = -70 -70 + 47 = -23 -23 = -23

OpenStudy (anonymous):

please someone help me with the question i made please